OpenStudy (anonymous):

Write the equation of a line parallel to the given line but passing through the given point. y = - x + 3; ( -2, -2) Write the equation of a line perpendicular to the given line but passing through the given point. y = - 2/3x + 2; (9, -3) Discuss the steps necessary to carry out each activity. Describe briefly what each line looks like in relation to the original given line. Answer these two questions briefly in your own words: 1. What does it mean for one line to be parallel to another? 2. What does it mean for one line to be perpendicular to another?

OpenStudy (shamim):

2 lines r parallel means their slopes r same

OpenStudy (shamim):

Can u tell me the slope of ur given line y=-x+3

OpenStudy (shamim):

U hv to compare with the standard equation of line y=mx+c

OpenStudy (shamim):

Here m=slope of line

OpenStudy (shamim):

A line which is going through a point x1,y1 is (y-y1)=m(x-x1)

OpenStudy (shamim):

Given x1=-2 y1=-2

OpenStudy (shamim):

So equation of a parallel line passing through -2,-2 is (y-(-2))=m(x-(-2)) Right?